Corp-X issued corporate bonds one year ago at par with a face value of $1000, an annual coupon rate of 6%(paid semi annually), and a 20 years to maturity.

At the moment, bonds of equivalent risk and maturity to these Corp-X bonds are being issued at par with a coupon rate of 5.5% per year(paid semi annually)

1. At the time that Corp-X bonds were issued, what was the Yield to Maturity of the bonds? And What is the current yield to Maturity of the Corp-X bonds?

2. What is the estimate of the current price of the Corp-X bond?

3. If you purchased the Corp-X bonds at the time they were issued, and held them over the past year, what is your realized return( IRR)? Express this as a stated annual interest rate, with semi-annual compounding.
Answer #1

1.
At the time of issue 6%
Current 5.5%

2.
=PV(5.5%/2,2*19,-6%*1000/2,-1000)=$1,058.48

3.
=RATE(2,30,-1000,1058.48)*2=11.6821%